A rectangle has a width of 3.5 inches and a length of x inches. If the diagonal of the rectangle is 12.5 inches, what is the value of x? Give the exact solution or round your answer to the nearest tenth
Answer by stanbon(75887)

A rectangle has a width of 3.5 inches and a length of x inches. If the diagonal of the rectangle is 12.5 inches, what is the value of x? Give the exact solution or round your answer to the nearest tenth
-----
Draw the picture.
You have a rectangle with width = 3.5 and length = x
The diagonal is the hypotenuse of of the triangle
formed by the width and the length.
Use Pythagoras to solve::
12.5^2 = x^2 + 3.5^2
x^2 = 144
x = 12 inches
------------------
Cheers,
Stan H.
--------------